Find the perimeter and area of a right triangle if one leg measures 11 yd. and the other leg measures 60 yd. Students must show all work step by step to receive full credit for the problem. Include correct units with your solution.
Answer by mananth(16946)

one leg measures 11 yd.
and the other leg measures 60 yd.
It is a right triangle
leg1^2+leg2^2= hypotenuse ^2
Hypotenuse = sqrt( 11 ^2+ 60 ^2)
= 61 m
Perimeter = 11+60+61=132 yards
...
Area = 1/2 * leg1 * leg2
Area= 0.5 * 11 * 60
Area = 330 yd^2
m.ananth@hotmail.ca